# Federated Continual Learning Goes Online: Uncertainty-Aware Memory Management for Vision Tasks and Beyond

### v3

@@ 1 Introduction
In recent years, federated learning (FL) (McMahan et al., 2017) has received increasing attention for its ability to allow local clients work towards the same objective without the need for sharing limited and sensitive information. However, despite the benefits provided by the privacy-preserving collaborative training, the assumptions of the standard scenario are far from realistic (Babakniya et al., 2024). For example, the standard static-single-task framework has very limited practical applicability in real-world cases where local clients often need to continuously learn new tasks. Let us consider the problem of classifying new COVID-19 variants as illustrative use-case. Given the evolving nature of the virus, new variants (i.e., new classes) arise over time. In this context, healthcare facilities which collaborate using FL would fail since the standard setting does not consider the dynamic increment of new classes. For this reason, a new paradigm was recently introduced to model more complex dynamics: federated continual learning (FCL) (Yoon et al., 2021a; Dong et al., 2022). This new paradigm takes the global-local communication and privacy-preserving abilities enabled by FL and combines them with the Continual Learning (CL) ability of learning different tasks sequentially over time. However, FCL not only shares the characteristics at the intersection of FL and CL, but also inherits their respective challenges. In particular one of the most prominent problems in CL, the so-called catastrophic forgetting for which the model is prone to suffer from significant performance degradation on older tasks. In CL, this problem is addressed in different ways ranging from memory-based approaches (Yoon et al., 2021b; Kumari et al., 2022; Hurtado et al., 2023) to generative methods (Shin et al., 2017; Lesort et al., 2019) or to regularization techniques (Lee et al., 2017; Aljundi et al., 2018; Chaudhry et al., 2018b; He and Jaeger, 2018). In the context of FCL, instead, the majority of the proposed approaches exploit generative models at both local and global level to trace and encode the past information and generate synthetic instances which faithfully mimic previous history whilst preserving data privacy. A substantial drawback of such generative approaches is that they are tailored to image data and it is not clear how they could translate to other data modalities.
Importantly, such generator-based approaches also imply an offline setting where task data are collected before training and remain unchanged over time: local generators require large training datasets of complete tasks and similarly, more recently proposed generative models based on data distillation can only be trained at the end of each task (Babakniya et al., 2024).
@@
Inspired by the problem of training models over online streams of data, we first formalise a new scenario to tackle the online problem in FCL (online-FCL). In this new scenario, we assume each client to learn from a stream of data where new data arrive in mini-batches which can only be processed once. As such, in line with the definition of online-CL, the model is updated with high frequency (Soutif-Cormerais et al., 2023). Then, taking inspiration from the most popular solutions in online-CL, we introduce memory buffers to alleviate CF at local level. More in detail, we propose an uncertainty-based memory management where data points are stored in local buffers according to their predictive uncertainty. Intuitively, predictive uncertainty provides a glimpse of the samples’ location in the decision space; samples with low uncertainty are the most representative ones for the respective class, while samples with high uncertainty represent data points that are close to the decision boundary and/or outliers. Thus, via estimates of predictive uncertainty, we can store in the memory samples with desired properties. Here, we propose to quantify uncertainty by directly estimating a generalized variance term from the loss function, which can be interpreted as a measure of epistemic uncertainty. To this end, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss (Gruber and Buettner, 2023) and estimate the Bregman Information (BI) as the variance term in logit space.
@@ 2 Related Work
Methods for class-IL can be divided into several categories (Mai et al., 2022) as follows; a) Regularization techniques alter the model parameter updates by adding penalty terms to the loss function (Lee et al., 2017; Aljundi et al., 2018), adjusting parameter gradients during optimization (Chaudhry et al., 2018b; He and Jaeger, 2018), or employing knowledge distillation (Rannen et al., 2017; Wu et al., 2019); b) Memory-based techniques exploit a fixed-size buffer containing samples from past tasks for replay (Aljundi et al., 2019; Chaudhry et al., 2019) or regularization (Nguyen et al., 2018; Tao et al., 2020); c) Generative-based techniques involve training generative models that can produce pseudo-samples mimicking data from the past (Shin et al., 2017; Lesort et al., 2019); d) Parameter-isolation-based techniques assign different model parameters to each task. This can be done either by activating only the relevant parameters for each task (Fixed Architecture) (Mallya and Lazebnik, 2018; Serra et al., 2018) or by adding new parameters and keeping the existing ones unchanged (Dynamic Architecture) (Aljundi et al., 2017; Yoon et al., 2018). In online-CL, where data arrive in single mini-batches and the model is updated frequently, rehearsal-based methods are favoured over more complex solutions, like generative methods, because they offer greater flexibility and require less training data and computational time (Mai et al., 2022).
@@ 3.2.1 Client-level
For each client , we introduce a fixed-size memory buffer . Similar to Chrysakis and Moens (2020), the memory population strategy is based on a class-balanced update, which is crucial to consider in case the stream of data is highly imbalanced. In fact, if classes are not equally represented in the memory, sampling from the memory may further deteriorate the predictive performance of the framework for under-represented classes. Differently from the populating strategy proposed in Chrysakis and Moens (2020), the criteria to decide which samples to keep in the memory is not random, but based on predictive uncertainty estimates in order to intentionally store samples with desired characteristics for each class. There are different ways to select the samples. We can decide to store a) the class-representative data points by selecting the least uncertain ones for each class (bottom-k); b) the easiest-to-forget samples by sampling the ones with high uncertainty (top-k).
@@
Different measure of predictive uncertainty can capture distinct aspects of a model’s irreducible aleatoric uncertainty (inherent in the data) and its epistemic uncertainty (that stems from the finite training data and can be reduced by gathering more data). In online CL, the most commonly used measures are derived directly from the confidence scores of the model and mostly capture the irreducible aleatoric uncertainty (Wimmer et al., 2023). Here, we hypothesize it may be more beneficial for the model to replay instances which are representative in the sense that there is low uncertainty about the data generating process (we refer to this as low epistemic uncertainty, while acknowledging that there exist varying definitions). To not rely on specialized Bayesian models, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss and compute a generalized variance term directly from the loss (Gruber and Buettner, 2023). Such bias-variance decomposition decomposes the expected prediction error (loss) into the model’s bias, variance, and an irreducible error (noise term). The latter is related to aleatoric uncertainty, whereas the variance term can directly be related to epistemic uncertainty (Gruber et al., 2023). This is the first attempt of using this decomposition for memory management in FCL contexts.
@@
Gruber and Buettner (2023) have recently shown that a bias-variance decomposition of cross-entropy loss gives rise to the Bregman Information as the variance term and measures the variability of the prediction in the logit space. We illustrate the different aspects of uncertainty captured by confidence scores and BI respectively in Figure 2. For example, data points close to the decision boundary have a low confidence score due to the inherently high aleatoric uncertainty; in contrast, due to the high density of observed data, there is actually a low uncertainty about the data generating process (DGP), resulting in a low BI (low epistemic uncertainty). Outliers far away from the decision boundary can have a high confidence score, but will have a high BI due to the high uncertainty regarding the DGP. We hypothesize that it is samples with a low BI that are most useful for replay in online-FCL.
To populate the memory with representative samples, we therefore propose using an uncertainty estimator based on Bregman Information (BI) (Gruber and Buettner, 2023). The authors demonstrate that BI at the sample level can be estimated through deep ensembles or test-time augmentation (TTA). However, to reduce the computational overhead at the local level, we employ TTA for computing the estimations. Let us consider the problem of multi-class classification (as in our case), where the standard loss is represented by the cross-entropy. Considering a set of perturbations for a given data point , we can compute the variance term of the classification loss as follows:
where represent the logit predictions and the LogSumExp (LSE) function respectively.
@@
Intuitively, a large value of means that the logits predicted across the perturbations vary significantly, suggesting a high uncertainty of the prediction and the DGP at this point of the input space. The use of this estimator is motivated by the fact that, in comparison with other uncertainty scores such as entropy, smallest margin, or least confidence, there is no information loss in the estimation step. In fact, if we inspect these alternative metrics (see Appendix A.1 for the equations), we can notice that they either need a normalization step to move the logits in the probability space or rest on the largest activation value only. Furthermore, as reported in Gruber and Buettner (2023); Ovadia et al. (2019), and Tomani and Buettner (2021), common confidence scores are reliable only in case of well-calibrated models.
In contrast, the BI-based estimation of the epistemic uncertainty is meaningful also under distribution shift and able to identify robust and representative samples (Gruber and Buettner, 2023).
@@ 4.1 Datasets and Settings
To understand the behaviour of different algorithms in the online-FCL scenario, we use two datasets commonly used in the literature, namely, CIFAR10 and CIFAR100 (Krizhevsky et al., 2009). We randomly assign a set of classes to 5 tasks and we split the task data evenly among the clients such that the task sequence for task and client does not share any data with the other clients. In this way, since the class assignment per task is different every time, we can identify the strategy providing greater flexibility and effectiveness irrespective of the composition of the tasks. Then, as anticipated in the introductory part, we assess the performance under more difficult and realistic conditions. Instead of using datasets that, to some extent, share similar characteristics with CIFAR (i.e., ImageNet) or represent unrealistic tasks (such as EMNIST), we validate our results on datasets for biomedical image analysis. Apart from the change in the domain (which in turn means different backgrounds in the images, different statistics, etc.), these datasets pose an additional challenge compared to the standard benchmark datasets, i.e., data imbalance. To reflect realistic conditions where recent tasks contain fewer data points than the older ones since the time to collect them is shorter, we assign classes to tasks based on the class size. We decide to focus on two biomedical datasets annotated by expert clinical pathologists; the colorectal cancer histology (CRC-Tissue) dataset (Kather et al., 2019; Yang et al., 2023) containing images divided in 8 classes of hematoxylin–eosin (HE)–stained slides taken from patients with colorectal cancer (CRC), and the kidney cortex cells (KC-Cell) dataset containing 8 classes of human kidney cortex tissue sections (Ljosa et al., 2012; Yang et al., 2023). Finally, to evaluate our approach on text classification tasks, we employ three different datasets: the 20NewsGroups dataset (Lang, 1995), DBPedia (Zhang et al., 2015), and Yahoo Answers (Zhang et al., 2015). As for vision tasks, we randomly assign classes to tasks every run. Statistics and more details about the used datasets can be found in the supplementary material (Appendix A.3).
@@
Given the novelty of the online setting, there are no direct competitors which we could refer to. We therefore investigate the most successful algorithms from the online-CL and the FCL literature. We compare our approach with standard baselines for FL, i.e., FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020), a standard memory-based approach for online-CL, namely Experience Replay (ER) (Chaudhry et al., 2019), and two state-of-the-art approaches for FCL with generative replay, namely MFCL (Babakniya et al., 2024) and FedCIL (Qi et al., 2023). Our decision to use ER is based on the fact that, despite its simplicity, it is surprisingly competitive when compared with more sophisticated and newer approaches as shown in recently conducted empirical surveys (Soutif-Cormerais et al., 2023). For FCL, we are particularly interested in MFCL because its data-free solution trains the generator on the server side. As such, we believe it may be potentially feasible in the online scenario, in contrast to other approaches such as FedCIL that train local generators at the client-side (Qi et al., 2023). In addition to the standard ER, we also include the class-balanced version (CBR) presented in Chrysakis and Moens (2020). Finally, to show the competitiveness of , we also consider other uncertainty scores. In particular, least confidence (LC), margin sampling (MS), ratio of confidence (RC), and entropy (EN) (Shannon, 1948; Campbell et al., 2000; Culotta and McCallum, 2005).
@@ 4.2 Empirical Results
From the experiments conducted on the CIFAR datasets, we can observe that the proposed approach is able to reduce CF consistently across different memory sizes and class-per-task assignments. The results in Tables 1 and 2 suggest that storing the class-representative samples (i.e., the least uncertain data points for each class) provides a benefit in terms of predictive performance gain and forgetting reduction. Our findings are substantiated in the classification tasks for biomedical images. Although the images originate from another domain, Table 3 confirms that in terms of CF, BI outperforms the considered memory-based baselines in most of the cases. Crucially, we also show the ability of our approach to work with imbalanced real-world data. Table 5 and Figure 3 summarise our findings; in comparison with both memory-based and generative-based methods, our simple-yet-elegant approach is able to perform consistently across tasks (left plot) and reduce CF on different datasets (right plot). Finally, we demonstrate that our method outperforms baselines also on non-vision data. Table 4 reports the results on the textual classification datasets. Here we find that the difference in terms of predictive accuracy is less marked. This is because the baselines have higher accuracy on the last tasks and poor performance on the first ones, while in our case the accuracy is kept more uniform across the tasks (reflected in a lower average forgetting) as shown in the left plot in Figure 3. Note that the use of generative baselines is restricted to image data only.

### v4

@@ 1 Introduction
In recent years, federated learning (FL) (McMahan et al., 2017) has received increasing attention for its ability to allow local clients work towards the same objective without the need for sharing limited and sensitive information. However, despite the benefits provided by the privacy-preserving collaborative training, the assumptions of the standard scenario are far from realistic (Babakniya et al., 2024). For example, the standard static-single-task framework has very limited practical applicability in real-world cases where local clients often need to continuously learn new tasks. Let us consider the problem of classifying new COVID-19 variants as illustrative use-case. Given the evolving nature of the virus, new variants (i.e., new classes) arise over time. In this context, healthcare facilities which collaborate using FL would fail since the standard setting does not consider the dynamic increment of new classes. For this reason, a new paradigm was recently introduced to model more complex dynamics: federated continual learning (FCL) (Yoon et al., 2021a; Dong et al., 2022). This new paradigm takes the global-local communication and privacy-preserving abilities enabled by FL and combines them with the Continual Learning (CL) ability of learning different tasks sequentially over time. However, FCL not only shares the characteristics at the intersection of FL and CL, but also inherits their respective challenges. In particular one of the most prominent problems in CL, the so-called catastrophic forgetting (CF) for which the model is prone to suffer from significant performance degradation on older tasks. In CL, this problem is addressed in different ways ranging from memory-based approaches (Yoon et al., 2021b; Kumari et al., 2022; Hurtado et al., 2023) to generative methods (Shin et al., 2017; Lesort et al., 2019) or to regularization techniques (Lee et al., 2017; Aljundi et al., 2018; Chaudhry et al., 2018b; He & Jaeger, 2018). In the context of FCL, instead, the majority of the proposed approaches exploit generative models at both local and global level to trace and encode the past information and generate synthetic instances which faithfully mimic previous history whilst preserving data privacy. A substantial drawback of such generative approaches is that they are tailored to image data and it is not clear how they could translate to other data modalities. Importantly, such generator-based approaches also imply an offline setting where task data are collected before training and remain unchanged over time: local generators require large training datasets of complete tasks and similarly, more recently proposed generative models based on data distillation can only be trained at the end of each task (Babakniya et al., 2024).
@@
Inspired by the problem of training models over online streams of data, we first formalise a new scenario to tackle the online problem in FCL (online-FCL). In this new scenario, we assume each client to learn from a stream of data where new data arrive in mini-batches which can only be processed once. As such, in line with the definition of online-CL, the model is updated with high frequency (Soutif-Cormerais et al., 2023). Then, taking inspiration from the most popular solutions in online-CL, we introduce memory buffers to alleviate CF at local level. More in detail, we propose an uncertainty-based memory management where data points are stored in local buffers according to their predictive uncertainty. Intuitively, predictive uncertainty provides a glimpse of the samples’ location in the decision space; samples with low uncertainty are the most representative ones for the respective class, while samples with high uncertainty represent data points that are close to the decision boundary and/or outliers. Thus, via estimates of predictive uncertainty, we can store in the memory samples with desired properties. Here, we propose to quantify uncertainty by directly estimating a generalized variance term from the loss function, which can be interpreted as a measure of epistemic uncertainty. To this end, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss (Gruber & Buettner, 2023) and estimate the Bregman Information (BI) as the variance term in logit space.
@@ 2 Related Work
Methods for class-IL can be divided into several categories (Mai et al., 2022) as follows; a) Regularization techniques alter the model parameter updates by adding penalty terms to the loss function (Lee et al., 2017; Aljundi et al., 2018), adjusting parameter gradients during optimization (Chaudhry et al., 2018b; He & Jaeger, 2018), or employing knowledge distillation (Rannen et al., 2017; Wu et al., 2019); b) Memory-based techniques exploit a fixed-size buffer containing samples from past tasks for replay (Aljundi et al., 2019; Chaudhry et al., 2019) or regularization (Nguyen et al., 2018; Tao et al., 2020); c) Generative-based techniques involve training generative models that can produce pseudo-samples mimicking data from the past (Shin et al., 2017; Lesort et al., 2019); d) Parameter-isolation-based techniques assign different model parameters to each task. This can be done either by activating only the relevant parameters for each task (Fixed Architecture) (Mallya & Lazebnik, 2018; Serra et al., 2018) or by adding new parameters and keeping the existing ones unchanged (Dynamic Architecture) (Aljundi et al., 2017; Yoon et al., 2018). In online-CL, where data arrive in single mini-batches and the model is updated frequently, rehearsal-based methods are favoured over more complex solutions, like generative methods, because they offer greater flexibility and require less training data and computational time (Mai et al., 2022).
@@ 3.2.1 Client-level
For each client , we introduce a fixed-size memory buffer . Similar to Chrysakis & Moens (2020), the memory population strategy is based on a class-balanced update, which is crucial to consider in case the stream of data is highly imbalanced. In fact, if classes are not equally represented in the memory, sampling from the memory may further deteriorate the predictive performance of the framework for under-represented classes. Differently from the populating strategy proposed in Chrysakis & Moens (2020), the criteria to decide which samples to keep in the memory is not random, but based on predictive uncertainty estimates in order to intentionally store samples with desired characteristics for each class. There are different ways to select the samples. We can decide to store a) the class-representative data points by selecting the least uncertain ones for each class (bottom-k); b) the easiest-to-forget samples by sampling the ones with high uncertainty (top-k).
@@
Different measure of predictive uncertainty can capture distinct aspects of a model’s irreducible aleatoric uncertainty (inherent in the data) and its epistemic uncertainty (that stems from the finite training data and can be reduced by gathering more data). In online CL, the most commonly used measures are derived directly from the confidence scores of the model and mostly capture the irreducible aleatoric uncertainty (Wimmer et al., 2023). Here, we hypothesize it may be more beneficial for the model to replay instances which are representative in the sense that there is low uncertainty about the data generating process (we refer to this as low epistemic uncertainty, while acknowledging that there exist varying definitions). To not rely on specialized Bayesian models, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss and compute a generalized variance term directly from the loss (Gruber & Buettner, 2023). Such bias-variance decomposition decomposes the expected prediction error (loss) into the model’s bias, variance, and an irreducible error (noise term). The latter is related to aleatoric uncertainty, whereas the variance term can directly be related to epistemic uncertainty (Gruber et al., 2023). This is the first attempt of using this decomposition for memory management in FCL contexts.
@@
Gruber & Buettner (2023) have recently shown that a bias-variance decomposition of cross-entropy loss gives rise to the Bregman Information as the variance term and measures the variability of the prediction in the logit space. We illustrate the different aspects of uncertainty captured by confidence scores and BI respectively in Figure 2. For example, data points close to the decision boundary have a low confidence score due to the inherently high aleatoric uncertainty; in contrast, due to the high density of observed data, there is actually a low uncertainty about the data generating process (DGP), resulting in a low BI (low epistemic uncertainty). Outliers far away from the decision boundary can have a high confidence score, but will have a high BI due to the high uncertainty regarding the DGP. We hypothesize that it is samples with a low BI that are most useful for replay in online-FCL. To populate the memory with representative samples, we therefore propose using an uncertainty estimator based on Bregman Information (BI) (Gruber & Buettner, 2023). The authors demonstrate that BI at the sample level can be estimated through deep ensembles or test-time augmentation (TTA). However, to reduce the computational overhead at the local level, we employ TTA for computing the estimations. Let us consider the problem of multi-class classification (as in our case), where the standard loss is represented by the cross-entropy. Considering a set of perturbations for a given data point , we can compute the variance term of the classification loss as follows:
@@
where represent the logit predictions and the LogSumExp (LSE) function respectively. Intuitively, a large value of means that the logits predicted across the perturbations vary significantly, suggesting a high uncertainty of the prediction and the DGP at this point of the input space. The use of this estimator is motivated by the fact that, in comparison with other uncertainty scores such as entropy, smallest margin, or least confidence, there is no information loss in the estimation step. In fact, if we inspect these alternative metrics (see Appendix A.1 for the equations), we can notice that they either need a normalization step to move the logits in the probability space or rest on the largest activation value only. Furthermore, as reported in Gruber & Buettner (2023); Ovadia et al. (2019), and Tomani & Buettner (2021), common confidence scores are reliable only in case of well-calibrated models. In contrast, the BI-based estimation of the epistemic uncertainty is meaningful also under distribution shift and able to identify robust and representative samples (Gruber & Buettner, 2023).
@@ 4.1 Datasets and Settings
To understand the behaviour of different algorithms in the online-FCL scenario, we use two datasets commonly used in the literature, namely, CIFAR10 and CIFAR100 (Krizhevsky et al., 2009), and TinyImageNet (Le & Yang, 2015). We randomly assign a set of classes to 5 tasks (10 for TinyImageNet) and we split the task data evenly among the clients such that the task sequence for task and client does not share any data with the other clients. In this way, since the class assignment per task is different every time, we can identify the strategy providing greater flexibility and effectiveness irrespective of the composition of the tasks. Then, as anticipated in the introductory part, we assess the performance under more difficult and realistic conditions. Instead of using datasets that, to some extent, share similar characteristics with CIFAR (i.e., ImageNet) or represent unrealistic tasks (such as EMNIST), we validate our results on datasets for biomedical image analysis. Apart from the change in the domain (which in turn means different backgrounds in the images, different statistics, etc.), these datasets pose an additional challenge compared to the standard benchmark datasets, i.e., data imbalance. To reflect realistic conditions where recent tasks contain fewer data points than the older ones since the time to collect them is shorter, we assign classes to tasks based on the class size. We decide to focus on two biomedical datasets annotated by expert clinical pathologists; the colorectal cancer histology (CRC-Tissue) dataset (Kather et al., 2019; Yang et al., 2023) containing images divided in 8 classes of hematoxylin–eosin (HE)–stained slides taken from patients with colorectal cancer (CRC), and the kidney cortex cells (KC-Cell) dataset containing 8 classes of human kidney cortex tissue sections (Ljosa et al., 2012; Yang et al., 2023). Finally, to evaluate our approach on text classification tasks, we employ three different datasets: the 20NewsGroups dataset (Lang, 1995), DBPedia (Zhang et al., 2015), and Yahoo Answers (Zhang et al., 2015). As for vision tasks, we randomly assign classes to tasks every run. Statistics and more details about the used datasets can be found in the supplementary material (Appendix A.3).
@@
Given the novelty of the online setting, there are no direct competitors which we could refer to. We therefore investigate the most successful algorithms from the online-CL and the FCL literature. We compare our approach with standard baselines for FL, i.e., FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020), a standard memory-based approach for online-CL, namely Experience Replay (ER) (Chaudhry et al., 2019), and two state-of-the-art approaches for FCL with generative replay, namely MFCL (Babakniya et al., 2024) and FedCIL (Qi et al., 2023). Our decision to use ER is based on the fact that, despite its simplicity, it is surprisingly competitive when compared with more sophisticated and newer approaches as shown in recently conducted empirical surveys (Soutif-Cormerais et al., 2023). For FCL, we are particularly interested in MFCL because its data-free solution trains the generator on the server side. As such, we believe it may be potentially feasible in the online scenario, in contrast to other approaches such as FedCIL that train local generators at the client-side (Qi et al., 2023). In addition to the standard ER, we also include the class-balanced version (CBR) presented in Chrysakis & Moens (2020). Finally, to show the competitiveness of , we also consider other uncertainty scores. In particular, least confidence (LC), margin sampling (MS), ratio of confidence (RC), and entropy (EN) (Shannon, 1948; Campbell et al., 2000; Culotta & McCallum, 2005).
@@ 4.2 Empirical Results
From the experiments conducted on the CIFAR datasets, we can observe that the proposed approach is able to reduce CF consistently across different memory sizes and class-per-task assignments. The results in Tables 1 and 2 suggest that storing the class-representative samples (i.e., the least uncertain data points for each class) provides a benefit in terms of predictive performance gain and forgetting reduction. Our findings are substantiated in the classification tasks for biomedical images. Although the images originate from another domain, Table 3 confirms that in terms of CF, BI outperforms the considered memory-based baselines in most of the cases. Crucially, we also show the ability of our approach to work with imbalanced real-world data. Table 5 and Figure 3 summarise our findings; in comparison with both memory-based and generative-based methods, our simple-yet-elegant approach is able to perform consistently across tasks (left plot) and reduce CF on different datasets (right plot). Finally, we demonstrate that our method outperforms baselines also on non-vision data. Table 4 reports the results on the textual classification datasets. Here we find that the difference in terms of predictive accuracy is less marked. This is because the baselines have higher accuracy on the last tasks and poor performance on the first ones, while in our case the accuracy is kept more uniform across the tasks (reflected in a lower average forgetting) as shown in the left plot in Figure 3. Note that the use of generative baselines is restricted to image data only. Results for TinyImageNet are reported in Appendix A.4 - Table 7.